# RGB-Event based Pedestrian Attribute Recognition: A Benchmark Dataset and An Asymmetric RWKV Fusion Framework

### v1

@@ 5.3 Comparison on Public Datasets
OTN-RWKV 87.66 84.78 89.03 89.38 89.07
@@
Threshold mA Acc Prec Recall F1
0.60 87.23 84.04 88.62 88.91 88.62
0.65 87.55 84.16 88.62 89.09 88.70
0.70 87.68 83.93 88.32 89.11 88.71
0.75 87.66 84.78 89.03 89.38 89.07
0.80 87.70 84.13 88.75 88.92 88.69
0.85 87.72 84.74 88.99 89.40 89.05
0.95 86.55 82.43 87.32 87.44 87.05
Table 8: Ablation study of different backbones and framework modules on the EventPAR dataset.
No. Backbone Aggregation Fusion EventPAR Dataset
Max Mean Gnn Similarity Concat Add 1x1Conv OTN-RWKV mA Acc F1
01 RWKV 87.56 84.62 88.98
02 RWKV 87.50 84.68 89.01
03 RWKV 84.53 82.24 87.50
04 RWKV 87.59 84.80 89.10
05 RWKV 87.50 84.68 89.01
06 RWKV 86.92 84.45 88.88
07 RWKV 87.13 84.40 88.86
08 RWKV 87.38 84.82 89.10
@@
11 RWKV 87.66 84.78 89.07
5.4 Ablation Study
In this section, we conduct a comprehensive ablation study on the core modules of EventPAR, including different input configurations, various backbone architectures, fusion strategies, event data aggregation methods, and the threshold settings in the similarity-based filtering mechanism.
$\bullet$ Analysis of different input settings. In this study, we varied the number of RGB inputs and conducted extensive experiments, as shown in Table 5. The first two rows represent RGB-only and Event-only inputs, both yielding strong results. RGB performed slightly worse due to added perturbations. When combining RGB and Event, performance improved, demonstrating their complementarity. However, excessive RGB inputs led to performance drops, likely due to redundant features and preprocessing noise. These factors increased the burden during feature extraction, introduced irrelevant cues, and weakened the ability to capture essential semantic information, ultimately degrading recognition performance.
$\bullet$ Analysis of different aggregation strategies of Event.  As shown in the first part of Table 8, we systematically compared four event data aggregation strategies: Max Pooling, Mean Pooling, a graph-based GNN, and our proposed Similarity-based approach. While Max and Mean pooling offer efficient but coarse feature representations, GNN captures structural dependencies with stronger expressive power. Our Similarity-based method adaptively fuses features based on semantic similarity between event frames, effectively retaining discriminative temporal information and filtering noise. n terms of mA 87.59, Accuracy 84.80, and F1 89.10, it outperformed all other strategies, hereby validating the effectiveness of our proposed similarity-based aggregation strategy.
$\bullet$ Analysis of different feature fusion methods.  We adopted two fusion strategies: holistic fusion and stepwise fusion. The corresponding results are shown in Table 6 and the second part of Table 8, respectively. We compared our proposed OTN-RWKV fusion strategy with three common methods: Concat, Add, and 1×1Conv. Although Concat increases feature dimensionality, it fails to capture inter-modal relationships. Add struggles with modality discrepancies, and 1×1Conv may lead to the loss of semantic details. In contrast, OTN-RWKV achieved superior performance in both fusion modes. It recorded mA of 87.66, ACC of 84.78, and F1 of 89.07 in holistic fusion, and mA of 87.38, ACC of 84.82, and F1 of 89.10 in stepwise fusion, outperforming all baseline methods. By finely integrating RGB and event modality features, the proposed strategy effectively captures their complementarity, significantly improves model robustness and overall performance, and fully validates its effectiveness.
$\bullet$ Analysis of different backbone.  To further evaluate the effectiveness of the proposed RWKV model, we conducted comparative experiments with ViT and ResNet-50 under identical settings. ViT captures global context, while ResNet-50 focuses on local features. In contrast, RWKV combines the strengths of RNNs and Transformers, achieving a balance between long-range dependency modeling and efficiency. Using consistent inputs and training protocols, our RWKV model outperformed both backbones with mA, Accuracy, and F1 of 87.66, 84.78, and 89.10, respectively, as shown in the third part of Table 8, highlighting its superiority in pedestrian attribute recognition.
@@ 5.4 Ablation Study
5.5 Visualization
$\bullet$ Visualization of Predicted Results.  As shown in Fig. 6, we present a visualization of attribute prediction results using different input modalities: RGB only, Event only, and the fused bimodal input. In the figure, green text represents the ground truth attributes, while red text indicates incorrectly predicted ones. The results show that using a single modality, whether the RGB data with added noise or the Event data alone, cannot achieve comprehensive and accurate predictions. In contrast, when both modalities are combined, the model can better capture complementary information, which significantly improves the accuracy and robustness of attribute recognition.
@@ 6 Conclusion and Future Works
In this paper, we addressed the limitations of existing pedestrian attribute recognition (PAR) methods, which are predominantly based on RGB frame cameras and suffer from challenges such as sensitivity to lighting conditions and motion blur. Moreover, current approaches focus primarily on external appearance and clothing attributes, neglecting the emotional dimensions of pedestrians. To overcome these constraints, we introduced a novel multi-modal RGB-Event attribute recognition task, leveraging the advantages of event cameras in low-light, high-speed, and low-power scenarios. Our key contribution is the creation of EventPAR, the first large-scale multi-modal PAR dataset, which includes 100K paired RGB-Event samples covering 50 attributes related to both appearance and six human emotions. By retraining and evaluating mainstream PAR models on EventPAR, we established a comprehensive benchmark, offering valuable data and algorithmic baselines for the community. Furthermore, we proposed a novel RWKV-based multi-modal PAR framework, incorporating an RWKV visual encoder and an asymmetric RWKV fusion module. Extensive experiments conducted on EventPAR and two simulated datasets demonstrated state-of-the-art performance, underscoring the effectiveness of our approach.
In our future works, we will focus on exploiting learning-based new event representations for the perception of event streams for pedestrian attribute recognition.

### v2

@@ 5.4 Ablation Study
OTN-RWKV 87.70 84.94 89.15 89.48 89.18
Table 7: Comparison of different aggregation strategies.
No. Max Mean GNN Sim mA Acc F1
01 87.56 84.62 88.98
02 87.50 84.68 89.01
03 84.53 82.24 87.50
04 87.59 84.80 89.10
$\bullet$ Analysis of different feature fusion methods.  As shown in Table 6, We compared our proposed OTN-RWKV fusion strategy with three common methods: Concat, Add, and 1×1Conv. OTN-RWKV achieved top results with mA 87.70, ACC 84.94, and F1 89.18, outperforming other methods. It effectively fuses RGB and event features, boosting robustness and performance.
$\bullet$ Analysis of different aggregation strategies of Event.  As shown in Table 7, we systematically compared four event data aggregation strategies: Max Pooling, Mean Pooling, a graph-based GNN, and our proposed Similarity-based approach. Our method adaptively fuses event features, preserves temporal cues, suppresses noise, and validates the effectiveness of the aggregation strategy.
@@
No. Threshold mA Acc Prec Recall F1
01 0.60 87.70 84.14 88.56 89.07 88.68
02 0.65 87.55 84.16 88.62 89.09 88.70
03 0.75 87.70 84.94 89.15 89.48 89.18
04 0.80 87.62 84.17 88.77 88.94 88.71
05 0.95 86.53 82.46 87.37 87.49 87.12
$\bullet$ Analysis of different backbone.  To validate the effectiveness of the RWKV, we conducted comparisons with ViT and ResNet-50 under the same settings. RWKV achieved superior performance with mA, Accuracy, and F1 of 87.70, 84.94, and 89.18, respectively, demonstrating its advantage in pedestrian attribute recognition, as shown in Table 9.
Table 9: Comparison of different backbones.
No. Backbone mA Acc F1
@@
03 RWKV 87.70 84.94 89.18
5.5 Visualization
$\bullet$ Visualization of Predicted Results.  As shown in Fig. 6, predictions using RGB, event, and fused bimodal inputs indicate that single modalities struggle to provide accurate and comprehensive recognition, while bimodal fusion better captures complementary information, significantly improving accuracy and robustness.
@@ 5.5 Visualization
$\bullet$ Visualization of emotion attribute prediction.  Fig. 7 visualizes multiple metrics for six emotion attributes. Our innovative RGB-Event fusion effectively captures emotional cues, consistently outperforming others and validating the value of emotion attributes in pedestrian recognition.
Figure 7: Visualization of emotion attribute prediction results by our method and other compared approaches.
@@ 6 Conclusion and Future Works
In this paper, we addressed the limitations of existing pedestrian attribute recognition (PAR) methods, which are predominantly based on RGB frame cameras and suffer from challenges such as sensitivity to lighting conditions and motion blur. Moreover, current approaches focus primarily on external appearance and clothing attributes, neglecting the emotional dimensions of pedestrians. To overcome these constraints, we introduced a novel multi-modal RGB-Event attribute recognition task, leveraging the advantages of event cameras in low-light, high-speed, and low-power scenarios. Our key contribution is the creation of EventPAR, the first large-scale multi-modal PAR dataset, which includes 100K paired RGB-Event samples covering 50 attributes related to both appearance and six human emotions. By retraining and evaluating mainstream PAR models on EventPAR, we established a comprehensive benchmark, offering valuable data and algorithmic baselines for the community. Furthermore, we proposed a novel RWKV-based multi-modal PAR framework, incorporating an RWKV visual encoder and an asymmetric RWKV fusion module. Extensive experiments conducted on EventPAR and two simulated datasets demonstrated state-of-the-art performance, underscoring the effectiveness of our approach. In our future works, we will focus on exploiting learning-based new event representations for the perception of event streams for pedestrian attribute recognition.